OpenStudy (anonymous):

a pouch contains 6 red, 5 white and 4 blue balls. the red balls are numbered from 1-6, white from 1-5, blue from 1-4. three balls are drawn in succession without replacement. what is the probabilty that the 3 balls drawn have the same number? answer key says it's 16/455 please explain. thanks!

OpenStudy (anonymous):

@Hero @ParthKohli @satellite73 @sauravshakya etc. etc? :D

OpenStudy (anonymous):

4*3/15 *2/14*1/13

OpenStudy (anonymous):

@sauravshakya 4*3/15 *2/14*1/13 =4/455 :( it's 16/455 according to the book O: :(

OpenStudy (anonymous):

I am pretty sure it is 4/455

OpenStudy (anonymous):

please explain why thanks!

OpenStudy (anonymous):

Probability of picking balls with number 1,1and 1 is 3/15*2/14*1/13

OpenStudy (anonymous):

Similarly probability of picking balls with number 2,2,2 :3,3,3:4,4,4 are 3/15*2/14*1/13

OpenStudy (anonymous):

So, probabilty that the 3 balls drawn have the same number = 3/15*2/14*1/13+3/15*2/14*1/13+3/15*2/14*1/13+3/15*2/14*1/13 =4*3/15*2/14*1/13 =4/455

OpenStudy (anonymous):

got it?